# A Cascaded Random Access Quantum Memory

### v1

@@ Abstract
Dynamic random access memory is critical to classical computing but notably absent in experimental quantum computers. Here we realize an 8-bit cascaded random access quantum memory using superconducting circuits and cavities and showcase the ability to perform arbitrary gate operations on it. In addition to individual error channels such as photon loss, quantum memories can also experience decoherence from many-body self-interaction. We characterize the origin and contributions of many-body infidelity throughout the memory cycle. We find that individual modes can be accessed with $\lesssim 1.5\%$ infidelity per mode and that the entire memory can be accessed in arbitrary order with an error rate below the depolarization threshold of the surface code, paving the way for fault-tolerant quantum memories.
@@ I Introduction
As we continue to demonstrate quantum error correction and make significant steps toward fault-tolerant quantum processors across multiple platforms [1, 2, 3, 4, 5, 6], a key challenge is scaling hardware resources to control many logical qubits efficiently. Classical computers tackle this challenge by assigning specialized roles for transistors. Although they consist of many transistors, only a small fraction perform logic operations. Most transistors are dedicated to random access memory (RAM), which allows energy-efficient control and access to large quantities of data. Similarly, a quantum computer can benefit from having separate components for logical processing and information storage. Compared to atomic systems where high coherent hyperfine states [7, 5] are naturally used as intrinsic memory, superconducting qubits– despite offering fast gates–lack dedicated storage subsystems in conventional architectures. Bridging this gap requires a resource that combines their fast logic with a coherent memory while allowing random access for efficient control: a random access quantum memory (RAQM).
Establishing a distinction between processing and storage components is also motivated by the underlying physics of various superconducting platforms. Nonlinear, strongly interacting systems like transmons [8] and fluxoniums [9] enable high-fidelity gates and are close to enabling fault-tolerant quantum computation. However, many promising data qubits are linear, weakly interacting systems, like photons in cavities [10, 11] and on-chip striplines [12]. Noise-protected superconducting quantum circuits [13, 14] are promising but often challenging to perform gates with. Arguably, the best memory qubit would be incapable of direct quantum interaction with other subsystems, either by symmetry or topology. RAQMs avoid forcing a single system to balance control and coherence, instead integrating optimized memories and processors to scale quantum systems efficiently [15].
Analogous to classical computing systems, RAQMs become truly advantageous when their storage capacity significantly exceeds that of their processing units, driving the need for scalable quantum memories. Apart from physically increasing memory components, quantum memory size can also be scaled in the frequency domain using a multimode system. These multimode memories have been realized using coplanar waveguides [16] and 3D microwave cavities [17, 18, 19]. This hardware-efficient approach enables random access of memory modes through a single channel. While larger RAQMs may be susceptible to idle errors, their extended coherence times and elevated tolerance to depolarization errors support scalable implementations. In fact, when integrated with appropriate fault-tolerant architectures, even modest RAQM sizes can result in multi-fold expansion in the capacity to control logical qubits.
Here, we introduce and realize a RAQM for superconducting qubits using 3D multimode cavities. As shown in Figure 1 (a-b), we employ a cascaded architecture that introduces a buffer layer of modes that suppresses direct interactions between the data and storage layers. As a result, the buffer layer protects the storage layer from inherited decoherence and stray interactions from the processor qubit. When scaled up as shown in Figure 1 (a), this architecture allows flexible encoding of logical qubits within the layers [20, 21] and enables transversal entangling gates between logical qubits within the same cavity set. The structure of the paper is as follows. We will first present the random access memory’s coherence, Hamiltonian, and protocols. Next, we benchmark each aspect of the RAQM’s operations, identifying both individual and many-body sources of decoherence.
II Random Access Quantum Memory Device
@@
Figure 1: (a) The proposed fault-tolerant architecture using Random Access Quantum Memory (RAQM), where each data qubit in the transmon processor (red) also serves as a register qubit for its corresponding RAQM. Each RAQM comprises a buffer mode (green) and a storage multimode (blue). Modes from different cavities are organized into layers to store logical qubits. Logical qubits are transferred from storage layers (blue) to the buffer layer for processing. (b) The RAQM device schematic diagram. The readout (black), buffer (green, $2$ modes), and storage (blue, $7$ modes) 3D flute cavities are fabricated on the same high-purity aluminum bulk. The transmon (red) is dispersively coupled to the readout and the buffer cavities, providing universal single buffer mode control. The RF-flux-modulated SQUID coupler (purple) enables parametrically actuated beam-splitter interactions between coupled cavity modes, allowing information transfer between the buffer and storage layers. (c) Frequency spectrum and mode coherence. The storage modes have frequencies evenly distributed between $5.3\ \mathrm{GHz}$ and $6.4\ \mathrm{GHz}$, with the highest $T_{1}$ exceeding $1.2\ \mathrm{ms}$. Error bars are smaller than marker sizes.
Our RAQM device (see Fig. 1 (b)) comprises two layers of information processing: the buffer (B) and storage (S) layers. Both layers are implemented using 3D superconducting flute cavities, which support multiple high-coherence bosonic modes [18]. These layers could also be implemented using on-chip [12], nanomechanical [22], or spin-based oscillators [23]. We realize swaps between the buffer and storage layers by flux pumping a SQUID coupler. The buffer modes are capacitively coupled to the transmon (Q), whose anharmonicity provides universal single-mode control. The storage modes serve as the quantum memory, accessed through multiplexed control of buffer modes: Any storage mode’s information is swapped into a buffer mode, modified through transmon control, and then swapped back into any storage mode. In a larger system, the transmon (or other type of processor) can be embedded as part of a quantum error correction circuit. In this work, information is encoded within the single-photon subspace $\{\left|0\right\rangle,\left|1\right\rangle\}$ of the modes, though the architecture is compatible with more complex bosonic [24] or dual-rail encodings [25, 26].
We begin with the Hamiltonian $H_{0}$ that describes the buffer and storage modes of the RAQM system:
$H_{0}$ $=\omega_{b}a_{b}^{{\dagger}}a_{b}+\sum_{j=1}^{7}\omega_{sj}a_{sj}^{{\dagger}}a_{sj}.$
Here, the annihilation operators and frequencies for the buffer mode $B$ and storage mode $S_{j}$ are denoted as ${a_{b},a_{sj}}$ and ${\omega_{b},\omega_{sj}}$. Although the buffer cavity contains two modes ($B_{1}$ and $B_{2}$), for our RAQM demonstration, we use only the lower frequency mode $B_{1}$ and thus remove its subscript for simplicity.
To transfer information between buffer and storage modes, we add a SQUID coupler [27] chip that provides programmable beamsplitter interactions $H_{\text{b.s.}}(t)$:
$H_{\text{b.s.}}(t)$ $=\sum_{i}\frac{g_{i}(t)}{2}\left(a_{b}^{{\dagger}}a_{si}+a_{b}a_{si}^{{\dagger}}\right).$
The coefficients $g_{i}(t)$ denote the beam-splitter interaction activated between buffer mode and storage $S_{i}$. To activate the beam-splitter interaction $g_{i}(t)$, the SQUID coupler is DC flux biased at $\Phi_{\text{DC}}=0.269\pi$ and RF modulated at the difference frequency between corresponding buffer and storage modes. This provides a beam-splitter rate proportional to the RF modulation amplitude. (See Supplementary Section B for detailed discussions).
To universally control the buffer mode, we add a transmon with frequency $\omega_{q}$ and anharmonicity $\alpha$ that is dispersively coupled to the buffer mode:
$H_{\chi_{qb}}$ $=\omega_{q}a_{q}^{\dagger}a_{q}+\frac{\alpha}{2}a_{q}^{\dagger}a_{q}^{\dagger}a_{q}a_{q}+\chi_{qb}a_{q}^{{\dagger}}a_{q}a_{b}^{{\dagger}}a_{b}.$
Here, the transmon annihilation operator is $a_{q}$, and the dispersive interaction between the transmon and $B_{i}$ is $\chi_{qbi}$.
The non-linearities present in the transmon and coupler introduce self-Kerr and cross-Kerr $\chi_{mn}$ interactions between buffer and storage modes:
$H_{\chi_{sb}}$ $=\sum_{m,n\in\{b,si\}}\left(\chi_{mn}a_{m}^{{\dagger}}a_{m}a_{n}^{{\dagger}}a_{n}\right).$
The qubit’s nonlinearity on the storage modes is largely suppressed by the cascaded architecture. The largest cross-Kerr strength between the buffer and storage modes is below $3.5\,\mathrm{kHz}$, while those between the storage modes are at the hundred-Hz level, inherited from the coupler.
The full RAQM system Hamiltonian $H$ is the sum of all components:
$H$ $=H_{0}+H_{\chi_{qb}}+H_{\chi_{sb}}+H_{\text{b.s.}}(t)$ (1)
For simplicity, we trace out the readout and coupler modes in the system Hamiltonian $H$. The storage cavity also contains $2$ lossy dump modes $D_{i}$. These dump modes are used for a fast, modular reset of the RAQM system to speed up all experiments (See Supplementary Section C for details). Detailed device parameters and circuits are provided in Supplementary Section A and Supplementary Section B.
@@
Figure 2: Single storage mode access. (a) Using the four-wave mixing interactions $\left|f0\right\rangle\leftrightarrow\left|g1\right\rangle$ between the transmon and the buffer mode, we can prepare $\left|1\right\rangle$ in the buffer in $0.615\,\mu\mathrm{s}$. (b) By modulating the coupler’s RF flux at the frequency difference between buffer and storage mode $S_{2}$, a $0.439\,\mathrm{MHz}$ beamsplitter interaction is activated. (c) Randomized Benchmarking (RB) of $BS_{2}$ swap. We prepare $\left|1\right\rangle$ in the buffer, execute RB sequences, and measure the state of buffer and storage (labeled as $\left|BS_{i}\right\rangle$). The top panel shows the raw and post-selected probability of measuring $\left|BS_{2}\right\rangle$ returning to $\left|10\right\rangle$. Post-selection keeps results with exactly one photon in the buffer-storage pair to benchmark non-decoherence errors. The RB pulse sequences are shown in (d). $\left|BS_{i}\right\rangle$ is extracted using two parity measurements and a swap. The transmon is actively reset to $\left|g\right\rangle$ between parity measurements. All $BS_{i}$ swap fidelities are shown in (e).
A significant technical challenge was developing an on-chip, DC-compatible, broadband, fast-flux line for 3D cavity systems. The flux line contains two parts: a bandstop filter and a balun. The flux line filter provides a $3$ GHz stop-band for both common and differential modes to protect the buffer, storage, and coupler modes. Additionally, we introduce an on-chip DC-compatible flux balun, which is a necessary component for smoothening the frequency passband and suppressing stop-band spikes caused by stray common modes (see Supplementary Section F for details on the fast-flux line design). Taken together, this flux design ensures high coherence times, up to $T_{1}=1.25$ ms and $T_{\phi}=1.5$ ms, for our storage modes (See Fig. 1(c) for modes’ coherence and frequencies measured at $\Phi_{\text{DC}}=0.269\pi$).
III Single Memory Access
@@
In this section, we will describe and benchmark each step of our RAQM’s information transfer process. Accessing each storage mode involves three steps. First, information is swapped from the storage to the buffer. After reading the memory, the buffer state interacts with the transmon processor, returning a modified state for memory writing. This is equivalent to performing a gate on the buffer mode. Finally, the modified buffer state is swapped back into the original storage mode.
@@
We prepare states, manipulate them, and perform readout of the buffer mode through the transmon. Their dispersive interaction enables a four-wave mixing interaction $\left|f0\right\rangle\leftrightarrow\left|g1\right\rangle$ to realize the transmon-buffer swap [28]. Together with $ge$ transmon gates, we can perform universal gates on the buffer mode in its $\{0,1\}$ subspace (See Supplementary Section H for detail). As shown in Figure 2(a), we prepare $\left|1\right\rangle$ in the buffer mode in $0.615\,\mu\mathrm{s}$ using this interaction. The transmon supports single qubit gates with a fidelity of $99.96\%$ and single-shot readout with a fidelity of $98.02\%$, both used for RAQM active reset and post-selection measurements.
@@
We realize buffer-storage swaps through the RF flux modulation of the SQUID coupler. To activate the beamsplitter interactions between any two modes, the external flux $\Phi_{\text{ext}}$ threading the SQUID loop is modulated at the difference frequency between the two modes with phase $\phi$: $\Phi_{\text{ext}}=\Phi_{\text{DC}}+\epsilon\cos(\omega_{d}t+\phi)$. Figure 2(b) shows the swap between the buffer mode and $S_{2}$ storage mode with a rate of $0.44\,\mathrm{MHz}$. In the present device, the swap rate is limited by the sideband frequency collisions between different storage-buffer pairs. Details about other buffer-storage swap rates are shown in the Supplementary Section B.
We characterize our buffer-storage swap gates using beam-splitter randomized benchmarking (RB) [27]. The RB gate set $\{X,Y,\pm X/2,\pm Y/2\}$ is implemented using our calibrated beam-splitter gate $U_{\text{b.s.}}(\phi)$. For example, we realize swap gates $X=(X/2)^{2}=U_{\text{b.s.}}^{2}(0)$ and $Y=(Y/2)^{2}=U_{\text{b.s.}}^{2}(\pi/2)$ by choosing different RF modulation phase $\phi$. The RB experiment starts by loading one photon in the buffer mode, and after a sequence of RB gates, we measure the populations of both modes. We use the transmon to perform two buffer parity measurements (See Fig. 2(d)), each consisting of two transmon $\pi_{ge}/2$ rotations separated by the waiting time $\pi/(\chi_{qb})$. Between two parity measurements, the transmon is actively reset to $\left|g\right\rangle$ via a conditional $\pi_{ge}$ pulse, and the storage and buffer state is also swapped before the second parity measurement. These parity measurements yield $P_{nm}$, the final-state probability of the buffer-storage pair being in $\left|nm\right\rangle(n,m\leq 1)$.
We reconstruct the storage-buffer state after RB by combining the measurement outcomes. Fig. 2(c) shows the RB result of $BS_{2}$ swap as an example. We extract both a raw $(P_{10}+P_{11})/(P_{00}+P_{01}+P_{10}+P_{11})$ and post-selected $P_{10}/(P_{10}+P_{01})$ process fidelity, allowing us to separately extract how much error is from photon loss. By fitting the process fidelity to exponential decay, we can extract the raw and post-selected swap fidelities, labeled as $F_{ri}$ and $F_{pi}$. For the $BS_{2}$ pair, we measured $F_{r2}=99.642\pm 0.002\%$ and $F_{p2}=99.963\pm 0.015\%$. All $BS_{i}$ swap infidelities are listed in Fig. 2(e). The swap infidelities after post-selection could be attributed to coupler heating and cavity dephasing. A detailed buffer-storage swap error budget is shown in Supplementary Section I.
IV Random Access Quantum Memory Performance

### v2

# arXiv:2503.13953v2[quant-ph] 03 Mar 2026

@@ Abstract
Dynamic random access memory (DRAM) is critical to classical computing but notably absent in current superconducting quantum processors. Integrating high-coherence memory units would enable resource-efficient control of logical qubits and allow the separate optimization of logic and storage subsystems. Here, we realize an 8-bit cascaded random access quantum memory (RAQM). By introducing a buffer layer between the processor and a multimode storage cavity, we leverage the control resources of a single transmon to address eight memory modes while isolating them from processor non-linearities. We demonstrate arbitrary random access with an average infidelity of $\lesssim 1.5\%$ per mode, characterizing the many-body interactions that dominate the error budget. This architecture enables a significant reduction in control lines per logical qubit and supports transversal operations within the memory module, establishing a scalable unit cell for fault-tolerant quantum architectures.
@@ I Introduction
As experimental demonstrations of quantum error correction (QEC) advance toward fault-tolerant operation [48, 58, 55, 56, 4, 1], the field faces a critical challenge: scaling the hardware resources required to control increasingly large numbers of logical qubits. Current architectures typically rely on a uniform grid of qubits where control lines scale linearly with qubit count. In classical computing, this scaling bottleneck is resolved by architectural specialization, with a small active processor and a large random access memory (RAM), separating logic from information storage to maximize density and efficiency. Similarly, realizing large-scale quantum processors will require dedicated, high-capacity quantum memories to complement logical processing units [2, 17, 51].
Developing such a memory is particularly urgent for superconducting circuits. Unlike atomic systems, which possess naturally high coherent hyperfine states that serve as intrinsic memories [65, 4], superconducting qubits lack a natural storage subsystem. The physical requirements for processing and storage are fundamentally conflicting: high-fidelity gates require strong nonlinearity and interaction (e.g. transmons  [31], fluxonium [42]), whereas long-lived storage favors linear, weakly interacting systems (e.g. superconducting cavities [53, 45, 49], nanomechanical resonators [40, 3, 15, 14], or spin ensembles [54]). Alternatively, noise protected qubits [21, 28, 22] promise long coherence in their protected state but may have slower gates than simpler qubits. A memory architecture resolves these conflicting demands by integrating long-lived memories with non-linear processor qubits, enabling the separate optimization of logic and coherence.
Here, we realize a “cascaded” random access quantum memory (RAQM) module for superconducting circuits. In this architecture, each processor qubit is integrated with a memory unit consisting of multiple storage modes. This approach allows the control resources of a single logical qubit to be multiplexed across $N$ memory layers, significantly expanding logical capacity without increasing the control line count (Fig. 1a). Crucially, this uses classical addressing to swap logical states between the storage layer and the processor and does not require coherent quantum addressing as in QRAM [20, 64, 43]. Beyond hardware efficiency, the all-to-all connectivity within an RAQM module enables transversal gates between memory layers, reducing the time overhead for multi-logical-qubit operations compared to edge-to-edge connectivity constraints [25, 68, 17, 2].
We implement a single unit cell of this architecture using a multimode 3D superconducting cavity system (Fig. 1b). We demonstrate universal control of this system, achieving a baseline swap infidelity of $0.5\%$ when accessing a single memory mode in isolation, a figure well below practical error-correction thresholds. Furthermore, we perform the first comprehensive benchmarking of the full RAQM module, characterizing the additional error overhead imposed by many-body decoherence between memory modes. We find that even in the presence of these interactions, the memory can be accessed in arbitrary order with an infidelity of $\lesssim 1.5\%$ per mode. Consequently, the aggregate error rate for memory sizes up to 7 modes remains below the depolarization threshold ($17\%$) of the surface code [66, 5, 13, 34], establishing the RAQM as a viable, scalable unit cell for fault-tolerant superconducting quantum computers.
II Assembling the Memories
@@
Figure 1: (a) Proposed fault-tolerant architecture using Random Access Quantum Memory (RAQM). Each data qubit in the transmon processor (red) also serves as a register qubit for its corresponding RAQM. Each RAQM consists of a buffer mode (green) and a storage multimode (blue). Modes from different cavities are organized into layers to store logical qubits. Logical qubits are transferred from storage layers (blue) to the buffer layer for processing. (b) RAQM device schematic. The readout (black), buffer (green, 22 modes), and storage (blue, 77 modes) 3D flute cavities are fabricated on the same high-purity aluminum bulk. The transmon (red) is dispersively coupled to the readout and buffer cavities, enabling universal single-buffer-mode control. The RF-flux-modulated SQUID coupler (purple) enables parametrically driven beam-splitter interactions between coupled cavity modes, transferring information between buffer and storage layers. (c) Frequency spectrum and mode coherence. The storage modes are evenly distributed between $5.3\,\mathrm{GHz}$ and $6.4\,\mathrm{GHz}$, with the highest $T_{1}$ exceeding $1.2\,\mathrm{ms}$ (see Supplementary Section I). Error bars are smaller than the marker size. Details about dump modes are provided in Supplementary Section II.
RAQM device comprises two layers (see Fig. 1(b)): a storage layer (S) and a buffer layer (B), with the buffer connected to a transmon processor for readout and gates. Both layers are implemented using 3D superconducting flute cavities, which are engineered to host multiple high-coherence bosonic modes within a single physical component [9]. Within the larger architecture of Fig. 1(a), this device would function as a single RAQM module: many long-lived modes in the storage layer are serviced by a common processor that is accessed through the buffer layer.
These layers have different functions inside the RAQM. The buffer layer serves as the processor-facing interface, engineered to strongly couple with the nonlinear processor while maintaining sufficient coherence. As such, it acts as a staging area where quantum states are temporarily stored for fast, programmable control. On the other hand, the storage layer is a bank of weakly coupled quantum memories optimized for coherence rather than gate operations. In the full RAQM architecture, these storage modes play the role of randomly addressable memory cells that can be selectively updated by the processor.
The nonlinear processor in our device is a transmon (Q) that is capacitively coupled to the buffer modes. Together, the transmon and a chosen buffer mode form a processing unit that enables universal single-mode control through optimal-control or sideband techniques [23, 18, 33, 26]. All state preparation, gates, and measurements in this work are implemented by driving the transmon and buffer, avoiding directly exposing the storage modes to non-linearity. Within the broader architecture, this transmon would belong to a large grid of qubits that periodically refreshes the memory through QEC.
To mediate communication between the buffer and storage layers, we introduce an RF-flux-modulated SQUID coupler [39]. By parametrically driving this coupler, we implement programmable beam-splitter interactions between the buffer mode and selected storage modes, enabling coherent swaps between a chosen storage cell and the buffer. This multiplexed control strategy enables a single processor to access many storage modes while keeping the memory layer largely idle and protected.
The physics of the combined processor-RAQM system is described by the Hamiltonian $H=H_{0}+H_{\text{proc}}+H_{\text{b.s.}}(t)+H_{\text{Kerr}}$. The fundamental harmonic resources are described by $H_{0}=\omega_{b}a_{b}^{\dagger}a_{b}+\sum_{j=1}^{7}\omega_{s_{j}}a_{s_{j}}^{\dagger}a_{s_{j}}$, where $a_{b}$ ($a_{s_{j}}$) and $\omega_{b}$ ($\omega_{s_{j}}$) denote the annihilation operator and frequency of the buffer (storage) mode $j$.
Universal control of these modes is enabled by the nonlinear transmon processor, which is dispersively coupled to the buffer mode:
$H_{\text{proc}}=\omega_{q}a_{q}^{\dagger}a_{q}+\frac{\alpha}{2}a_{q}^{\dagger}a_{q}^{\dagger}a_{q}a_{q}+\chi_{qb}a_{q}^{\dagger}a_{q}a_{b}^{\dagger}a_{b},$ (1)
where $\omega_{q}$, $\alpha$, and $\chi_{qb}$ are the transmon frequency, anharmonicity, and dispersive shift, respectively. To enable memory access, the SQUID coupler is modulated to generate time-dependent beam-splitter interactions:
$H_{\text{b.s.}}(t)=\sum_{j}\frac{g_{j}(t)}{2}\left(a_{b}^{\dagger}a_{s_{j}}+a_{b}a_{s_{j}}^{\dagger}\right).$ (2)
Here, $g_{j}(t)$ represents the tunable coupling rate. Finally, residual nonlinearities from the transmon and coupler introduce self- and cross-Kerr interactions $H_{\text{Kerr}}=\sum_{m,n}\chi_{mn}a_{m}^{\dagger}a_{m}a_{n}^{\dagger}a_{n}$, where indices run over the set $\{b,s_{1},\dots,s_{7}\}$. Due to the cascaded architecture, the largest buffer-storage cross-Kerr is below $3.5\,\mathrm{kHz}$, while storage-storage terms remain at the hundred-Hz level.
In this work, we encode quantum information in the single-photon subspace $\{\left|0\right\rangle,\left|1\right\rangle\}$ of the cavity modes. The same hardware is also compatible with more complex bosonic encodings [44] and dual-rail architectures [59, 35], in which the buffer–storage connectivity enables hardware-efficient error correction and logical operations.
@@
Figure 2: Single storage mode access. (a) Using four-wave mixing interactions $\left|f0\right\rangle\leftrightarrow\left|g1\right\rangle$ between the transmon and buffer mode, we prepare $\left|1\right\rangle$ in the buffer within $0.615\,\mu\mathrm{s}$. (b) By modulating the coupler’s RF flux at the frequency difference between the buffer and storage mode $S_{2}$, a $0.439\,\mathrm{MHz}$ beam-splitter interaction is activated. (c) Randomized Benchmarking (RB) of the $BS_{2}$ swap. We prepare $\left|1\right\rangle$ in the buffer, execute RB sequences, and measure the joint buffer–storage state $\left|BS_{i}\right\rangle$. The top panel shows raw and post-selected probabilities of measuring $\left|BS_{2}\right\rangle$ returning to $\left|10\right\rangle$. Post-selection keeps outcomes with exactly one photon in the buffer–storage pair, isolating non-decoherence errors. The RB pulse sequences are shown in (d). $\left|BS_{i}\right\rangle$ is extracted using two parity measurements and a swap, with the transmon actively reset to $\left|g\right\rangle$ between parity checks. Swap fidelities for all $BS_{i}$ are shown in (e). RAQM reset protocols are discussed in Supplementary Section III.
III Controlling Single Quantum Memories
@@
In this section, we will describe and benchmark each step of our RAQM’s information transfer process. Accessing a specific memory mode proceeds in three steps: (1) a read operation swaps the quantum state from the target storage mode $S_{i}$ into the buffer $B$; (2) a gate operation modifies the state in the buffer via the transmon processor; and (3) a write operation swaps the modified state back into $S_{i}$. We characterize each stage of this cycle below.
@@
We prepare states, manipulate them, and perform readout of the buffer mode through the transmon. Their dispersive interaction enables a four-wave mixing interaction $\left|f0\right\rangle\leftrightarrow\left|g1\right\rangle$ to realize the transmon-buffer swap [50]. Together with geg-e transmon gates, we can perform universal gates on the buffer mode in its $\{0,1\}$ subspace (See Supplementary Section VIII for details). As shown in Fig. 2(a), we prepare $\left|1\right\rangle$ in the buffer mode in $0.615\,\mu\mathrm{s}$ using this interaction. The transmon supports single-qubit gates with a fidelity of $99.96\%$ and single-shot readout with a fidelity of $98.02\%$. To measure the buffer mode population, we use the transmon to perform the parity measurement, consisting of two transmon $\pi_{ge}/2$ rotations separated by a waiting time of $\pi/(\chi_{qb})$.
@@
We realize buffer-storage swaps through the RF flux modulation of the SQUID coupler. To activate the beamsplitter interactions between any two modes, the external flux $\Phi_{\text{ext}}$ threading the SQUID loop is driven as $\Phi_{\text{ext}}=\Phi_{\text{DC}}+\epsilon\cos(\omega_{d}t+\phi)$, with DC bias $\Phi_{DC}=0.269\pi$ and modulation frequency $\omega_{d}$. By choosing the modulation phase ($\phi$) and amplitude ($\epsilon$) appropriately, we obtain the calibrated beamsplitter gate $U_{\text{b.s.}}(\phi)$. Fig. 2(b) shows the swap between the buffer mode and $S_{2}$ storage mode with a rate of $0.44\,\mathrm{MHz}$. A key innovation of this work is the on-chip, broadband fast-flux line designed for 3D-cavity systems. It preserves cavity mode coherence–achieving $T_{1}$ up to $1.25\,\mathrm{ms}$ (see Fig. 1(c))–while enabling fast beamsplitter interaction rates of $0.3\text{--}0.5\,\mathrm{MHz}$ (see Supplementary Section VI for details on the fast-flux-line design). In the present device, the swap rate is limited by the sideband frequency collisions between different storage-buffer pairs. Details about other buffer-storage swap rates are shown in the Supplementary Section II.
We quantify swap performance using beam-splitter randomized benchmarking (RB) [39]. The sequence (Fig. 2(d)) initializes a single photon, applies random beam-splitter gates, and measures the joint parity. We define raw fidelity based on the recovery of the ideal state, and a post-selected fidelity by conditioning on the preservation of total photon number, $P_{10}/(P_{10}+P_{01})$, where $P_{nm}$ is the population of a buffer-storage pair having $n$ and $m$ photons. As shown in Fig. 2(e), we achieve an average raw (post-selected) fidelity of $99.32\%$ ($99.71\%$) across all $B-S_{i}$ pairs. The remaining incoherent error is attributed to coupler heating and cavity dephasing (see Supplementary Section X). While the current swap time ($\sim 0.5\,\mu\mathrm{s}$) is limited by frequency crowding, faster exchange rates are compatible with this architecture. Crucially, the measured swap infidelity satisfies even the stringent $\sim 1\%$ threshold typically required for active two-qubit gates in the surface code [1]. Since memory access operations occur less frequently than active gates, this performance provides a substantial margin for fault-tolerant operation.
IV Benchmarking Random Access Quantum Memory